Mal87:

Moxie wants to have $5000. How much money does she have to deposit in an account at 6% interest, compounded 3 times per year, in order to have $5000 at the end of 4 years? A. $4002.73 B. $3807.90 C. $3942.47 D. $4372.49

Shadow:

\[P = P_{i} \times (1 + \frac{ r }{ K })^{TK} \] Where Pi = the initial amount, r = the rate k = the of times compounded per year t = the # of years \[5000 = P_{i} \times (1 + \frac{ 0.06 }{ 3 })^{3 \times 4} \] \[5000 = P_{i} \times (1.02)^{12} \] \[\frac{ 5000 }{ 1.02^{12} }= P_{i} \] \[P_{i} = 3942.47\]

Shadow:

Basically just inputting variables then solving for Pi

Shadow:

All you really need to know is the formula and what each variable represents, then any problem is relatively easy.